Exercise · ZW-BERNOULLI-DISTRIBUTION-005 · ★☆☆ Basic

A fair die is rolled and \(X=1\) if a multiple of 3 shows, \(X=0\) otherwise. Compute \(E[X]\).

Choose one option.

Show answer
  • A. \(\tfrac13\)

Explanation

The multiples of 3 are 3 and 6, so \(p=\tfrac26=\tfrac13\) and \(E[X]=p\).

Study Probability and statistics in ZWMath →

Review the concept: Bernoulli distribution

A Bernoulli variable \(X\) only takes the values 1 (success) and 0 (failure); its only parameter \(p\) is the probability \(P\) of success.

\[P(X=1)=p,\quad P(X=0)=1-p\]

More practice problems: bernoulli distribution

← ZW-BERNOULLI-DISTRIBUTION-004 ZW-BERNOULLI-DISTRIBUTION-006 →